# An Image is Worth 1/2 Tokens After Layer 2: Plug-and-Play Inference Acceleration for Large Vision-Language Models

### v2

@@ 5.2 Model Settings
We test FastV with various open source models. For image understanding tasks, we conduct experiments on LLaVA1.5-7B, 13B [^1] https://github.com/haotian-liu/LLaVA Liu et al. (2023b), and Qwen-VL [^2] https://github.com/QwenLM/Qwen-VL Bai et al. (2023). When it comes to video understanding tasks, our baseline model is VideoLLaVA [^3] https://github.com/PKU-YuanGroup/Video-LLaVA Lin et al. (2023a). We adopt the settings as reported in their paper for the baseline models.
@@ 5.3 Main Results
Balance between Budget and Performance.
Considering the parameters of FastV, a lower $K$ and higher $R$ would generally result in a more significant FLOPs reduction with relative performance drop. Above from the decreased inference budget, it is interesting to note that sometimes FastV could lead to better performance than vanilla decoding. This bolsters our theory that FastV can be seen as a redundancy reduction operation.
@@
(d) w/ FastV (K=2, R=50%, on system prompt) 89.2 64.3 69.2 33.8
(e) w/ FastV (K=2, Prune first half system prompt) 17.5 27.8 Failed Failed
(f) w/ FastV (K=2, R=50%, on instruction) 77.3 50.1 56.5 29.5
@@ 5.4 Ablation Studies
We conduct ablation studies to evaluate the impact of various components within FastV, with the findings detailed in Figure 7 and Table 5.
Influence of Different K-R Combinations on Performance.
@@
In our previous observation about attention efficiency, we find out that the system prompt takes up of most attention even if they carry the least semantic information in the context. We conduct another experiment by directly prune the first half tokens of the system prompt. Comparing line (d) and (e), we can find that the head tokens in the system prompt have dominant effect on the model performance. Our findings also align with StreamingLLM Xiao et al. (2023) where they find that the first 4 tokens in LLM play the most important role during inference.
Compare with LLM’s Effective Attention Method.
As we find that LVLM and LLM bear similar attention patterns that the head tokens play an important role during inference, we wonder that whether the same attention optimization technique curated for LLM can be transfered to LVLM. Following the methodology outlined in StreamingLLM Xiao et al. (2023), we implemented an attention optimization scheme that incorporates an attention sink for the first four tokens and keep all local attention (context size=4) for the preceding four tokens throughout the decoding process. The experiment’s outcome, presented in line (g) of Table 5, indicates a substantial degradation in LVLM’s performance when this specific attention pattern is applied. This suggests a fundamental difference in how image tokens, as opposed to text tokens, contribute to the information processing within LLMs. This observation highlights the necessity for further investigation into the unique behaviors of image tokens within the LVLM framework, suggesting that the direct application of LLM attention/KV-Cache optimization mechanisms to LVLMs requires adaptation to accommodate the distinct characteristics of visual information.
5.5 Limitations
The FLOPs reduction ratio is based on the theoretical calculation considering the removal of image tokens, while actual inference budget can be influenced by a variety of factors such as inference framework optimization, specific CUDA kernels and hardwares. We are working on integrating FastV into mainstream LLM inference frameworks such as vLLM Kwon et al. (2023) for broader application.
@@ 6 Conclusion
In this paper, we propose FastV, a plug-and-play inference budget optimization method for Large Vision-Language Models. Our insight for FastV arises from our observation that the attention computation over visual tokens is of extreme inefficiency in the deep layers of popular LVLMs though they take up a large portion of input tokens. FastV prunes out the unnecessary visual tokens according to the attention score ranking, which results in significant inference budget reduction without sacrificing performance.
Appendix A Prompts for Different Tasks
Image Caption.
For image captioning tasks Nocaps and Flickr30k, we adopt prompt as “Describe the image in one sentence.”
AOKVQA.
For AOKVQA, we adopt the the multiple choice version of evaluation and use prompt as: “Analyse the image and choose the best answer for the following question: {question} Options: {options}. Output the letter of the correct answer.”
MMMU and PCA-Bench.
For evaluation in MMMU and PCA-Bench, we use the default prompts for each example as provided in the official dataset https://huggingface.co/datasets/MMMU/MMMU and https://huggingface.co/datasets/PCA-Bench/PCA-Bench-V1.
OCRVQA.
For OCRVQA, we use the default question as prompt for each example as provided in the official dataset.
Video QA.
For all video QA tasks, we use the default question as the prompt as provided in Video-LLaVA [^4] https://github.com/PKU-YuanGroup/Video-LLaVA/blob/main/TRAIN_AND_VALIDATE.md, and use the same tool from Video-ChatGPT [^5] https://github.com/mbzuai-oryx/Video-ChatGPT/blob/main/video_chatgpt/eval/run_inference_benchmark_general.py to conduct GPT evaluation.

### v3

@@ 5.2 Model Settings
We test FastV with various open source models. For image understanding tasks, we conduct experiments on LLaVA1.5-7B, 13B Liu et al. (2023b), and Qwen-VL Bai et al. (2023). When it comes to video understanding tasks, our baseline model is VideoLLaVA Lin et al. (2023). We adopt the settings as reported in their paper for the baseline models.
@@ 5.3 Main Results
Fine-grained Benchmarks and More Models
We conduct additional experiments with InstructBLIP and also with more fine-grained LVLM benchmarks such as SciQA-IMGLu et al. (2022), SeedBench Li et al. (2023a), MMVet Yu et al. (2023), and MME Fu et al. (2023), together with benchmarks requiring more visual processing such as AI2Diagram. The results and fine-grained scores of MME are shown in Table 2 and Table 3. FastV works well on different LVLM benchmarks with competitive performance. We find that InstructBLIP shows slightly more performance degradation than LLaVA with same FastV config. The gap soon closes when we just set K to 5. We think it’s because Q-Former initially reduces image tokens, resulting in direct information loss. Consequently, it requires adjusting the FastV parameters to avoid too much information loss.
@@
(d) w/ FastV (system prompt) 89.2 64.3 69.2 33.8
(e) w/ FastV (prune first half system prompt) 17.5 27.8 Failed Failed
(f) w/ FastV (instruction) 77.3 50.1 56.5 29.5
@@ 5.4 Ablation Studies
Balance between Cost and Performance.
@@
In our previous observation about attention efficiency, we find out that the system prompt takes up of most attention even if they carry the least semantic information in the context. We conduct another experiment by directly prune the first half tokens of the system prompt. Comparing line (d) and (e), we can find that the head tokens in the system prompt have dominant effect on the model performance. Our findings also align with StreamingLLM Xiao et al. (2023) where they find that the first 4 tokens in LLM play the most important role during inference. However, direcly applying the same sparse attention pattern as StreamingLLM would lead to a substantial degradation in LVLM’s performance as shown in line (g) of Table 7. This suggests a fundamental difference in how image tokens, as opposed to text tokens, contribute to the information processing within LLMs.
@@ 6 Conclusion
In this paper, we propose FastV, a plug-and-play inference cost optimization method for Large Vision-Language Models. Our insight for FastV arises from our observation that the attention computation over visual tokens is of extreme inefficiency in the deep layers of popular LVLMs though they take up a large portion of input tokens. FastV prunes out the unnecessary visual tokens according to the attention score ranking, which results in significant inference cost reduction without sacrificing performance.
Appendix A Appendix
Figure 8: Full Attention Maps of Each Layer of LLaVA.
Appendix B Evaluation Tasks Description
Image Captioning.
Image captioning requires the model to generate a description for a given image. We choose Nocaps Agrawal et al. (2019) and Flickr30k Plummer et al. (2015) as benchmarks and report CIDEr score Vedantam et al. (2015) as metric. For image captioning tasks Nocaps and Flickr30k, we adopt prompt as “Describe the image in one sentence.”
Visual Question Answering (VQA).
VQA requires the model to generate an answer for a given image-question pair. We select the development set of A-OKVQA Schwenk et al. (2022) and the test set of OCR-VQA  Mishra et al. (2019) as the benchmark and the report the multiple choice (MC) score of AOKVQA and Rouge-L score of OCR-VQA. For AOKVQA, we adopt the the multiple choice version of evaluation and use prompt as: “Analyse the image and choose the best answer for the following question: {question} Options: {options}. Output the letter of the correct answer.” For OCRVQA, we use the default question as prompt for each example as provided in the official dataset.
Multimodal Reasoning.
Compared with VQA, multimodal reasoning requires more advanced perception, knowledge and reasoning skills of the model, which are more suitable benchmarks to evaluate the integrated abilities of LVLMs. We choose MMMU and PCA-Bench Chen et al. (2024) as benchmarks. MMMU is a multimodal benchmark featuring multi-discipline tasks demanding college-level subject knowledge and reasoning skills. PCA-Bench is a complex embodied reasoning benchmark with error localization, which features three different domains including autonomous driving, robot and game. We report the multiple choice accuracy for the development set of MMMU and Perception, Cognition, Action, Genuine PCA scores for both the open and closed test set of PCA-Bench. We use the default prompts for each example as provided in the official dataset MMMU and PCA-Bench.
Video Question Answering.
Similar to VQA for single image, Video Question Answering requires the model to generate answer given a video-question pair. Current LVLMs usually deal with video question answering tasks by sampling multiple frames as input, resulting in longer image token sequences. We choose TGIF-QA Jang et al. (2017), MSVD-QA Xu et al. (2017b) and MSRVTT-QA Xu et al. (2017a) as benchmarks following the evaluation pipeline of Video-ChatGPT Maaz et al. (2023) and report the accuracy and chatgpt-score as metrics. We use the first 1K examples in each benchmark in our experiments due to the limited commercial API usage in evaluation. For all video QA tasks, we use the default question as the prompt as provided in Video-LLaVA, and use the same tool from Video-ChatGPT to conduct GPT evaluation.
Fine-grained Benchmarks
For the evaluation of the influence of FastV on LVLM performance, we incorporate four distinct Fine-grained benchmarks: MME Fu et al. (2023), Seed-Bench Li et al. (2023a), SciQA-IMG Lu et al. (2022), and MMVet Yu et al. (2023). MME offers a comprehensive evaluation of models’ perception and cognition abilities across a diverse set of tasks, focusing on intuitive and quantifiable analysis without extensive prompt engineering. SEED-Bench, on the other hand, evaluates generative comprehension across multiple dimensions, ensuring question relevance and quality through a mix of automated filtering and manual verification. While MME and SEED-Bench cover general abilities of LVLMs, SciQA-IMG and MMVet focus on the advanced aspects of multi-modal understanding. SciQA-IMG is a large-scale multimodal science question dataset annotated with detailed lectures and explanations. MMVet evaluates LVLMs on complex multimodal tasks, emphasizing multi-modal understanding and free-form answering capabilities, thus offering a comprehensive view of model performance.